MIT-BIH Arrhythmia Database Directory

Third Edition 23 July 1992

Harvard-MIT Division of Health Sciences and Technology Biomedical Engineering Center Copyright \begin{matrix}c \\ ◯ \end{matrix} 1980, 1988, 1992 Massachusetts Institute of Technology.

For information on the content or format of the database described in this directory, for information on related databases and software, or to obtain a copy of the database, write to:

MIT-BIH Database Distribution

Massachusetts Institute of Technology 77 Massachusetts Avenue, Room 20A-113 Cambridge, MA 02139 USA

Additional printed copies of this directory are available for $25 each (including shipping by surface mail) by writing to the address above. (Please make checks payable to “Beth Israel Hospital Biomedical Engineering Division.”)

Permission is granted to make and distribute verbatim copies of this directory provided that the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this directory under the conditions for verbatim copying, provided also that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this directory into another language, under the above conditions for modified versions.